OpenStudy (anonymous):

Help please . 1/3(-2^2-12) I got -10.68 but i dont know it its right.

jimthompson5910 (jim_thompson5910):

(1/3)(-2^2-12) (1/3)(-4-12) (1/3)(-16) -16/3 ... which is the exact answer -5.3333 ... which is the approximate answer

OpenStudy (anonymous):

noooo thats not right because you have to do the distributive property

jimthompson5910 (jim_thompson5910):

The original problem is \[\Large \frac{1}{3}\left(-2^2 - 12\right)\] correct?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1}{3}\left(-2^2 - 12\right)\] \[\Large \frac{1}{3}\left(-4 - 12\right)\] \[\Large \frac{1}{3}(-4)+\frac{1}{3}(-12)\] \[\Large -\frac{4}{3}-\frac{12}{3}\] \[\Large \frac{-4-12}{3}\] \[\Large -\frac{16}{3}\]

jimthompson5910 (jim_thompson5910):

So we get the same answer

OpenStudy (anonymous):

ok thank you.

jimthompson5910 (jim_thompson5910):

np